# 3DSRBench: A Comprehensive 3D Spatial Reasoning Benchmark

### v3

@@ Abstract
3D spatial reasoning is the ability to analyze and interpret the positions, orientations, and spatial relationships of objects within the 3D space. This allows models to develop a comprehensive understanding of the 3D scene, enabling their applicability to a broader range of areas, such as autonomous navigation, robotics, and AR/VR. While large multi-modal models (LMMs) have achieved remarkable progress in a wide range of image and video understanding tasks, their capabilities to perform 3D spatial reasoning on diverse natural images are less studied. In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, with 2,772 manually annotated visual question-answer pairs across 12 question types. We conduct robust and thorough evaluation of 3D spatial reasoning capabilities by balancing the data distribution and adopting a novel FlipEval strategy. To further study the robustness of 3D spatial reasoning w.r.t. camera 3D viewpoints, our 3DSRBench includes two subsets with 3D spatial reasoning questions on paired images with common and uncommon viewpoints. We benchmark a wide range of open-sourced and proprietary LMMs, uncovering their limitations in various aspects of 3D awareness, such as height, orientation, location, and multi-object reasoning, as well as their degraded performance on images with uncommon camera viewpoints. Our 3DSRBench provide valuable findings and insights about future development of LMMs with strong 3D spatial reasoning abilities. Our project page and dataset is available here.
@@ 1 Introduction
Recent large multi-modal models (LMMs) [1, 48, 4] have achieved significant improvements in a wide range of image and video understanding tasks, such as image captioning [2, 35], visual question answering [28, 24, 39, 13, 52], visual grounding [58], decision making [42, 33, 10], and action recognition [57, 43]. Notably, the spatial reasoning ability [30, 28, 17, 54], i.e., parsing 2D and 3D spatial relationships between objects, serves as a crucial foundation for various high-level reasoning and interaction in downstream tasks. Studying the spatial reasoning ability of current LMMs will help us identify specific types of factual errors, uncover their fundamental limitations, and inform targeted improvements to further advance current LMMs.
Prior datasets [30, 28, 36, 31] studying spatial relationships often focused on relationships w.r.t. the viewer, e.g., object A is to the left of object B from the viewer’s perspective. We regard these as 2D spatial relationships as they can be captured merely from 2D bounding boxes of the objects (see Fig. 2(b)). They neglect 3D spatial relationships in the 3D world space or those from an object’s perspective. Capturing 3D spatial relationships between objects in the images would help LMMs understand and predict the interactions between objects, and enable a broader range of applications in 3D, e.g., robotics and embodied AI.
Figure 1: Overview of our 3DSRBench. (a) Example questions from the four main types of 3D spatial reasoning questions, i.e., height, location, orientation, and multi-object reasoning. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy to remove left/right biases in 3D with paired VQAs (see Sec. 3.4).
To study how LMMs can capture 3D spatial relationships, previous works often exploited synthetic environments and generated images with 3D ground-truths [55, 56]. Visual question-answer pairs were automatically synthesized by applying pre-defined rules to the known 3D scene graphs and object attributes. The synthetic images exhibit a significant domain gap with natural images and lacked the diversity and richness in real-world. More recent works [17] explored real datasets with 3D annotations, e.g., Omni3D [9]. However, images in these datasets are limited to specific domains, such as indoor rooms and self-driving scenes. In general, visual question-answer pairs generated with rule-based methods from 3D annotations (i) limit the scope of theirs datasets to a small set of rigid object categories, and (ii) cannot enable a fine-grained and robust evaluation of 3D spatial relationships that can only be achieved with human annotated datasets (see Sec. 3.1).
@@
Another challenge of 3D spatial reasoning arises from the 6D viewpoint of the camera, i.e., the 3D location and 3D orientation from which we are viewing the 3D scene. As shown in Fig. 3, 3D spatial reasoning questions can be easier for common 6D viewpoints, e.g., ones positioned at the eye level with natural viewing angles, while being more challenging for other uncommon viewpoints. Although uncommon viewpoints are less populated in most image datasets, cameras in embodied AI and robotics are often positioned in these uncommon viewpoints. Hence it is of crucial importance for LMMs to retain good 3D spatial reasoning performance for both common and uncommon viewpoints. To fairly compare the 3D spatial reasoning capabilities of LMMs w.r.t. different camera viewpoints, we annotate another 672 visual question-answer pairs on multi-view synthetic images rendered from the HSSD dataset [32].
@@
To enable a comprehensive and robust evaluation of 3D spatial reasoning abilities, 3DSRBench adopts several key designs: (1) balanced data distributions in multiple aspects, such as balanced answer distribution and complementary images pairs that lead to opposite answers given the same question (see Fig. 1(b)); (2) avoiding questions with shortcuts or trivial answers; and (3) a novel FlipEval strategy for robust evaluation of 3D spatial reasoning abilities.
@@ 2 Related Works
In this section we introduce 3DSRBench for comprehensively analyzing the 3D spatial reasoning capabilities of LMMs. We start by presenting the design considerations in Sec. 3.1, i.e., how these design choices lead to a robust and valuable evaluation of 3D spatial reasoning capabilities. Then we show the four main question types in Sec. 3.2, as well as the challenges in each type of questions. Next we introduce the three splits of 3DSRBench and their scopes in Sec. 3.3. In Sec. 3.4 we present our evaluation strategies, including CircularEval and FlipEval. Please refer to Sec. A in supplementary materials where we provide details of our data collection and summary statistics of 3DSRBench.
@@ 3.1 Design of 3DSRBench
When developing 3DSRBench, we incorporate the following four key designs to enable a robust and valuable evaluation of 3D spatial reasoning capabilities. First, our 3D spatial reasoning questions are based on open-vocabulary entities. Previous spatial reasoning benchmarks [12, 52] largely relied on existing datasets with 3D annotations [9], which limited their scope to a small number of rigid object categories. In our 3DSRBench, we annotate 3D spatial reasoning questions across a broad range of open-vocabulary entities (see Fig. 1), enabling a thorough analysis of the 3D awareness and 3D reasoning capabilities of LMMs over diverse, commonly encountered real-world objects. Next, we avoid questions with shortcuts or trivial answers. For instance, objects higher in 3D space are usually higher in 2D space. We collect diverse VQAs and avoid those with clear shortcuts (see Fig. 2(a)). Also, when comparing which of the two objects has a smaller 3D distance to a third anchor object, we avoid the cases when there is a significant gap between the two distances, which lead to trivial answers. Moreover, we implement a balanced data distribution in various aspects, such as a roughly same number of yes/no answers and complementary image pairs [24] that lead to opposite answers given the same 3D spatial reasoning question (see Fig. 1(b)). This effectively removes priors in the answer distribution, e.g., pedestrians are often located lower than street lights, or the fact that objects higher in 3D space are also higher in 2D image plane. This design ensures that models cannot exploit biases or shortcuts for a higher benchmark performance. Lastly, we adopt special evaluation strategies for robust evaluation, including previous CircularEval [41] and our novel FlipEval (see Sec. 3.4).
@@ 3.2 Question Types
We present the 4 types of 3D spatial reasoning questions in our 3DSRBench. We discuss why they are challenging for LMMs and what kinds of 3D awareness and 3D spatial reasoning are needed to succeed in each type of questions. We present an overview of the 4 question types in Tab. 1.
@@
There are three subtypes of location-related questions, i.e., determining (i) if two objects are next to or far from each other, (ii) which of the two objects is closer to the camera, and (iii) if an object is directly above or underneath another object. Models must not only ground the 2D locations of the objects, but also understand the depth of field presented in the image. Consider the location question in Fig. 1(a). Although the 2D locations of the man and the hydrant are close, they are in fact far away from each other in the 3D space. Humans can determine the answer by estimating a rough depths of the two objects, or from other visual cues, such as how the pedestrian walk leads towards the vanishing point. Other examples include the top two questions in Fig. 1(b), which also require an understanding of the depth field.
@@
Orientation-related questions study the 3D spatial reasoning that involves estimating the 3D orientation of an object. These questions are divided into three subtypes: determining which “side” of an object faces the camera, whether an object is in front of or behind another, and if an object is positioned on the left or right side of another. Unlike previous 2D spatial reasoning questions [12] that focus on spatial relationships w.r.t. the viewer’s perspective, our orientation-related questions emphasize spatial relationships from the object’s perspective. As demonstrated in Fig. 2(b), 2D spatial reasoning questions can be addressed by analyzing objects’ 2D locations and depths. Meanwhile, our orientation questions require estimating objects’ 3D orientation and perform 3D spatial reasoning across various dimensions of 3D information.
@@
Figure 2: Challenges of 3D spatial reasoning questions in our 3DSRBench. See Sec. 3.2. (a) Height questions requires 3D spatial reasoning over a combination of camera extrinsics and object 3D locations. Notice how different camera pitch rotations play a crucial role to determine the final answer. (b) Previous 2D spatial reasoning questions can be addressed by analyzing objects’ 2D locations and depths, while our orientation questions require complex 3D spatial reasoning on objects’ 3D orientations and 3D locations.
@@ 3.3 Benchmark Splits
With the HSSD 3D scenes and controllable photorealistic rendering, we obtain multi-view images of the same 3D scene, each rendered with a common and an uncommon viewpoint. We ask the same 3D spatial reasoning question regarding the two images and study if models can obtain the correct answers on common and uncommon camera 6D viewpoints. We define “common” viewpoints as 6D camera poses with zero roll rotation, small pitch rotation, and taken from the height of a human, simulating the typical perspective when people take pictures. Conversely, “uncommon” viewpoints include 6D poses with noticeable roll rotation, large pitch rotation, or perspectives taken close to the ground or from a high location. The two synthetic splits are denoted by synthetic-common and synthetic-uncommon and examples from the two splits are demonstrated in Fig. 3. Notice how the answers by GPT-4o are correct when shown the image from a common camera 6D viewpoint and wrong when prompted from an uncommon viewpoint, despite both images present a clear view of the 3D scene and humans can derive the correct answers without any difficulty.
@@
Figure 3: Examples of the paired visual question-answer data in our 3DSRBench-synthetic. (a) Example questions from the four main types of 3D spatial reasoning questions. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy (see Sec. 3.4).
@@ 3.4 Evaluation
We further propose FlipEval to remove left/right biases in 3D with paired visual question-answer pairs. By applying horizontal flip to the image, we obtain a new visual question. The answer would generally remain the same, such as for location and height questions, but when it involves 3D spatial relationships such as “left” and “right”, the answer would change. We illustrate this idea in Fig. 1(b), where the elephant logo is the on the left of the truck but changes to the right side after horizontal flipping. FlipEval effectively removes left/right biases in 3D spatial relationship, such as driver often sitting on the left side of the car or most people holding tools in their right hands. Lastly FlipEval also helps to avoid the influence of random guessing and enriches the image distribution in our 3DSRBench.
@@ 4 Experiments
We first introduce our experimental settings in Sec. 4.1. Next in Sec. 4.2 we benchmark various LMMs on 3DSRBench. We further study how various model designs, i.e., choice of visual encoders and scaling of language models, attribute to the 3D spatial reasoning abilities. Then we evaluate various LMMs on our 3DSRBench-synthetic and analyze the robustness of LMMs w.r.t. uncommon camera viewpoints in Sec. 4.3. Lastly we present some failure cases of GPT-4o and Gemini 2.0 in Sec. 4.4, highlighting limitations of current state-of-the-art LMMs and discussing possible future improvements.
@@ 4.2 Results on 3D Spatial Reasoning Abilities
LLaVA-v1.5-7B [38] 38.1 39.1 46.9 28.7 34.7
Cambrian-1-8B [52] 42.2 23.2 53.9 35.9 41.9
LLaVA-NeXT-8B [40] 48.4 50.6 59.9 36.1 43.4
InternVL2.5-8B [16] 50.9 45.9 68.1 38.7 43.3
QWen2.5-VL-7B [7] 48.4 44.1 62.7 40.6 40.5
@@
Claude-3.5V-Sonnet [4] 48.2 53.5 63.1 31.4 41.3
Gemini-2.0-Flash [23] 49.8 49.7 68.9 32.2 41.5
Gemini-2.0-Flash-bbox [23] 47.5 45.2 66.5 27.7 41.4
Gemini-2.0-Flash-think [23] 51.1 53.0 67.1 35.8 43.6
GPT-4o-mini [29] 39.7 44.3 52.4 21.0 36.5
GPT-4o [29] 44.2 53.2 59.6 21.6 39.0
QWenVLMax [7] 52.0 45.1 70.7 37.7 44.8
Table 2: Experimental comparison of state-of-the-art large multi-modal models on our 3DSRBench. Results show that state-of-the-art LMMs exhibit limited 3D spatial reasoning capabilities. Please refer to Sec. 4.2 for detailed analyses.
@@
Table 4: Experimental results on our 3DSRBench-synthetic-common and 3DSRBench-synthetic-uncommon. We study the robustness of 3D spatial reasoning capabilities of LMMs by analyzing the performance gap between the two splits with images from the same 3D scene but from “common” and “uncommon” viewpoints. We find that LMMs does not generalize well to images with 6D camera viewpoints less represented in their training set. See Sec. 4.3 for detailed discussions.
We benchmark a wide range of open-sourced and proprietary LMMs on our 3DSRBench real and analyze the 3D spatial reasoning abilities on different types of questions. We consider three baseline results: (i) random: a simple baseline that predicts random answers for all visual questions. (ii) random++: a stronger random baseline that predicts consistent answers given different choice orders of a same visual question in CircularEval. (iii) human: a human-level performance established by human evaluators that did not participate in the data annotation process. We report the full results in Tab. 2.
@@
We study how design choices of visual encoders can benefit 3D spatial reasoning abilities. Built on LLaVA-v1.5-7B [39], we experiment on a range of models with different choices of visual foundation models, i.e., CLIP [47], MAE [25], DINOv2 [46], SAM [34], or model designs, i.e., mixed encoders and visual projectors. Results in Tab. 3 show that with mixed encoders, DINOv2 can improve the overall 3D spatial reasoning abilities of LMMs, specifically for orientation and multi-object reasoning questions that build heavily on object 3D orientations. We also notice significant improvements for height questions when adopting MAE and SAM as vision encoder, suggesting that having richer visual features could help localize objects better. With spatial vision aggregator (SVA) [52], we can further improve the LMM with mixed encoder from 37.2% to 37.8%, demonstrating that fusing the semantic features with 3D-aware features from DINOv2 would benefit subsequent reasoning.
@@
We study how the scaling of language model, i.e., in terms of the number of parameters, can help improve the 3D spatial reasoning abilities of LMMs. We consider two series of open-sourced LMMs, QWen2.5 [7] and InternVL2.5 [16], with a range of language model sizes from 0.5B to 72B. From the results shown in Fig. 4, we see that the scaling of language model sizes can effectively improve the 3D spatial reasoning abilities of LMMs. Larger language models with more parameters exhibit enhanced reasoning abilities. They can better capture 3D-aware information from the visual features and perform more complicated 3D spatial reasoning. However, given the importance of 3D spatial reasoning in a broad range of downstream applications, scaling up language model size is highly inefficient — LMMs with over 70B parameters exceed the computation capacity of common robotics or embodied AI systems and significantly limit the model throughput.

### v4

@@ Abstract
3D spatial reasoning is the ability to analyze and interpret the positions, orientations, and spatial relationships of objects within the 3D space. This allows models to develop a comprehensive understanding of the 3D scene, enabling their applicability to a broader range of areas, such as autonomous navigation, robotics, and AR/VR. While large multi-modal models (LMMs) have achieved remarkable progress in a wide range of image and video understanding tasks, their capabilities to perform 3D spatial reasoning on diverse natural images are less studied. In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, with 2,772 manually annotated visual question-answer pairs across 12 question types. We conduct robust and thorough evaluation of 3D spatial reasoning abilities by balancing data distribution and adopting a novel FlipEval strategy. To further study the robustness of 3D spatial reasoning w.r.t. camera 3D viewpoints, our 3DSRBench includes two subsets with 3D spatial reasoning questions on paired images with common and uncommon viewpoints. We benchmark a wide range of open-sourced and proprietary LMMs, uncovering their limitations in various aspects of 3D awareness, such as height, orientation, location, and multi-object reasoning, as well as their degraded performance on images from uncommon 6D viewpoints. Our 3DSRBench provide valuable findings and insights about future development of LMMs with strong spatial reasoning abilities. Our project page is available here.
@@ 1 Introduction
Recent large multi-modal models (LMMs) [1, 50, 4] have achieved significant improvements in a wide range of image and video understanding tasks, such as image captioning [2, 34], visual question answering [27, 23, 38, 54], visual grounding [60], decision making [41, 32, 10], and action recognition [59, 42]. Notably, the spatial reasoning ability [29, 27, 16, 56], i.e., parsing 2D and 3D spatial relationships between objects, serves as a crucial foundation for various high-level reasoning and interaction in downstream tasks. Studying the spatial reasoning ability of current LMMs will help us identify specific types of factual errors, uncover their fundamental limitations, and inform targeted improvements to further advance current LMMs.
Prior datasets [29, 27, 35, 30] studying spatial relationships often focused on relationships w.r.t. the viewer, e.g., object A is to the left of object B from the viewer’s perspective. We regard these as 2D spatial relationships as they can be captured merely from 2D bounding boxes of the objects (see Fig.˜2(b)). They neglect 3D spatial relationships in the 3D world space or those from an object’s perspective. Capturing 3D spatial relationships between objects in the images would help LMMs understand and predict the interactions between objects, and enable a broader range of applications in 3D, e.g., robotics and embodied AI.
Figure 1: Overview of our 3DSRBench. (a) Example questions from the four main types of 3D spatial reasoning questions, i.e., height, location, orientation, and multi-object reasoning. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy to remove left/right biases in 3D with paired VQAs (see Sec.˜3.4).
To study how LMMs can capture 3D spatial relationships, previous works often exploited synthetic environments and generated images with 3D ground-truths [57, 58]. Visual question-answer pairs were automatically synthesized by applying pre-defined rules to the known 3D scene graphs and object attributes. The synthetic images exhibit a significant domain gap with natural images and lacked the diversity and richness in real-world. More recent works [16] explored real datasets with 3D annotations, e.g., Omni3D [9]. However, images in these datasets are limited to specific domains, such as indoor rooms and self-driving scenes. In general, visual question-answer pairs generated with rule-based methods from 3D annotations (i) limit the scope of theirs datasets to a small set of rigid object categories, and (ii) cannot enable a fine-grained and robust evaluation of 3D spatial relationships that can only be achieved with human annotated datasets (see Sec.˜3.1).
@@
Another challenge of 3D spatial reasoning arises from the 6D viewpoint of the camera, i.e., the 3D location and 3D orientation from which we are viewing the 3D scene. As shown in Fig.˜3, 3D spatial reasoning questions can be easier for common 6D viewpoints, e.g., ones positioned at the eye level with natural viewing angles, while being more challenging for other uncommon viewpoints. Although uncommon viewpoints are less populated in most image datasets, cameras in embodied AI and robotics are often positioned in these uncommon viewpoints. Hence it is of crucial importance for LMMs to retain good 3D spatial reasoning performance for both common and uncommon viewpoints. To fairly compare the 3D spatial reasoning capabilities of LMMs w.r.t. different camera viewpoints, we annotate another 672 visual question-answer pairs on multi-view synthetic images rendered from the HSSD dataset [31].
@@
To enable a comprehensive and robust evaluation of 3D spatial reasoning abilities, 3DSRBench adopts several key designs: (1) balanced data distributions in multiple aspects, such as balanced answer distribution and complementary images pairs that lead to opposite answers given the same question (see Fig.˜1(b)); (2) avoiding questions with shortcuts or trivial answers; and (3) a novel FlipEval strategy for robust evaluation of 3D spatial reasoning abilities.
@@ 2 Related Works
In this section we introduce 3DSRBench for comprehensively analyzing the 3D spatial reasoning capabilities of LMMs. We start by presenting the design considerations in Sec.˜3.1, i.e., how these design choices lead to a robust and valuable evaluation of 3D spatial reasoning capabilities. Then we show the four main question types in Sec.˜3.2, as well as the challenges in each type of questions. Next we introduce the three splits of 3DSRBench and their scopes in Sec.˜3.3. In Sec.˜3.4 we present our evaluation strategies, including CircularEval and FlipEval. Please refer to Sec.˜A in supplementary materials where we provide details of our data collection and summary statistics of 3DSRBench.
@@ 3.1 Design of 3DSRBench
When developing 3DSRBench, we incorporate the following four key designs to enable a robust and valuable evaluation of 3D spatial reasoning capabilities. First, our 3D spatial reasoning questions are based on open-vocabulary entities. Previous spatial reasoning benchmarks [12, 54] largely relied on existing datasets with 3D annotations [9], which limited their scope to a small number of rigid object categories. In our 3DSRBench, we annotate 3D spatial reasoning questions across a broad range of open-vocabulary entities (see Fig.˜1), enabling a thorough analysis of the 3D awareness and 3D reasoning capabilities of LMMs over diverse, commonly encountered real-world objects. Next, we avoid questions with shortcuts or trivial answers. For instance, objects higher in 3D space are usually higher in 2D space. We collect diverse VQAs and avoid those with clear shortcuts (see Fig.˜2(a)). Also, when comparing which of the two objects has a smaller 3D distance to a third anchor object, we avoid the cases when there is a significant gap between the two distances, which lead to trivial answers. Moreover, we implement a balanced data distribution in various aspects, such as a roughly same number of yes/no answers and complementary image pairs [23] that lead to opposite answers given the same 3D spatial reasoning question (see Fig.˜1(b)). This effectively removes priors in the answer distribution, e.g., pedestrians are often located lower than street lights, or the fact that objects higher in 3D space are also higher in 2D image plane. This design ensures that models cannot exploit biases or shortcuts for a higher benchmark performance. Lastly, we adopt special evaluation strategies for robust evaluation, including previous CircularEval [40] and our novel FlipEval (see Sec.˜3.4).
@@ 3.2 Question Types
We present the 4 types of 3D spatial reasoning questions in our 3DSRBench. We discuss why they are challenging for LMMs and what kinds of 3D awareness and 3D spatial reasoning are needed to succeed in each type of questions. We present an overview of the 4 question types in Tab.˜1.
@@
There are three subtypes of location-related questions, i.e., determining (i) if two objects are next to or far from each other, (ii) which of the two objects is closer to the camera, and (iii) if an object is directly above or underneath another object. Models must not only ground the 2D locations of the objects, but also understand the depth of field presented in the image. Consider the location question in Fig.˜1(a). Although the 2D locations of the man and the hydrant are close, they are in fact far away from each other in the 3D space. Humans can determine the answer by estimating a rough depths of the two objects, or from other visual cues, such as how the pedestrian walk leads towards the vanishing point. Other examples include the top two questions in Fig.˜1(b), which also require an understanding of the depth field.
@@
Orientation-related questions study the 3D spatial reasoning that involves estimating the 3D orientation of an object. These questions are divided into three subtypes: determining which “side” of an object faces the camera, whether an object is in front of or behind another, and if an object is positioned on the left or right side of another. Unlike previous 2D spatial reasoning questions [12] that focus on spatial relationships w.r.t. the viewer’s perspective, our orientation-related questions emphasize spatial relationships from the object’s perspective. As demonstrated in Fig.˜2(b), 2D spatial reasoning questions can be addressed by analyzing objects’ 2D locations and depths. Meanwhile, our orientation questions require estimating objects’ 3D orientation and perform 3D spatial reasoning across various dimensions of 3D information.
@@
Figure 2: Challenges of 3D spatial reasoning questions in our 3DSRBench. See Sec.˜3.2. (a) Height questions requires 3D spatial reasoning over a combination of camera extrinsics and object 3D locations. Notice how different camera pitch rotations play a crucial role to determine the final answer. (b) Previous 2D spatial reasoning questions can be addressed by analyzing objects’ 2D locations and depths, while our orientation questions require complex 3D spatial reasoning on objects’ 3D orientations and 3D locations.
@@ 3.3 Benchmark Splits
With the HSSD 3D scenes and controllable photorealistic rendering, we obtain multi-view images of the same 3D scene, each rendered with a common and an uncommon viewpoint. We ask the same 3D spatial reasoning question regarding the two images and study if models can obtain the correct answers on common and uncommon camera 6D viewpoints. We define “common” viewpoints as 6D camera poses with zero roll rotation, small pitch rotation, and taken from the height of a human, simulating the typical perspective when people take pictures. Conversely, “uncommon” viewpoints include 6D poses with noticeable roll rotation, large pitch rotation, or perspectives taken close to the ground or from a high location. The two synthetic splits are denoted by synthetic-common and synthetic-uncommon and examples from the two splits are demonstrated in Fig.˜3. Notice how the answers by GPT-4o are correct when shown the image from a common camera 6D viewpoint and wrong when prompted from an uncommon viewpoint, despite both images present a clear view of the 3D scene and humans can derive the correct answers without any difficulty.
@@
Figure 3: Examples of the paired visual question-answer data in our 3DSRBench-synthetic. (a) Example questions from the four main types of 3D spatial reasoning questions. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy (see Sec.˜3.4).
@@ 3.4 Evaluation
Following the left-right biases discussed in [49], we further propose a novel FlipEval to remove left/right biases in 3D with paired visual question-answer pairs. By applying horizontal flip to an image, we obtain a new visual question. The answer would generally remain the same, such as for location and height questions, but when it involves 3D spatial relationships such as “left” and “right”, the answer would change. We illustrate this idea in Fig.˜1(b), where the elephant logo is the on the left of the truck but changes to the right after image flipping. FlipEval effectively removes left/right biases in 3D spatial relationship, such as driver often sitting on the left side of the car or most people holding tools in their right hands. Lastly FlipEval also avoids the influence of random guessing and enriches the image distribution in our 3DSRBench.
@@ 4 Experiments
We first introduce our experimental settings in Sec.˜4.1. Next in Sec.˜4.2 we benchmark various LMMs on 3DSRBench. We further study how various model designs, i.e., choice of visual encoders and scaling of language models, attribute to the 3D spatial reasoning abilities. Then we evaluate various LMMs on our 3DSRBench-synthetic and analyze the robustness of LMMs w.r.t. uncommon camera viewpoints in Sec.˜4.3. Lastly we present some failure cases of GPT-4o and Gemini 2.0 in Sec.˜4.4, highlighting limitations of current state-of-the-art LMMs and discussing possible future improvements.
@@ 4.2 Results on 3D Spatial Reasoning Abilities
LLaVA-v1.5-7B [37] 13 38.1 39.1 46.9 28.7 34.7
Cambrian-1-8B [54] 11 42.2 23.2 53.9 35.9 41.9
LLaVA-NeXT-8B [39] 6 48.4 50.6 59.9 36.1 43.4
InternVL2.5-8B [15] 4 50.9 45.9 68.1 38.7 43.3
QWen2.5-VL-7B [7] 6 48.4 44.1 62.7 40.6 40.5
Specialist
SpatialLLM [45] 9 44.8 45.8 61.6 30.0 36.7
SpatialRGPT [16] 14 32.7 55.9 39.0 27.8 20.0
SpatialRGPT w/ depth [16] 6 48.4 55.9 60.0 34.2 42.3
SpatialReasoner [44] 1 60.3 52.5 75.2 55.2 51.8
@@
Claude-3.5V-Sonnet [4] 7 48.2 53.5 63.1 31.4 41.3
Gemini-2.0-Flash [22] 5 49.8 49.7 68.9 32.2 41.5
Gemini-2.0-Flash-bbox [22] 8 47.5 45.2 66.5 27.7 41.4
Gemini-2.0-Flash-think [22] 3 51.1 53.0 67.1 35.8 43.6
GPT-4o-mini [28] 12 39.7 44.3 52.4 21.0 36.5
GPT-4o [28] 10 44.2 53.2 59.6 21.6 39.0
QWenVLMax [7] 2 52.0 45.1 70.7 37.7 44.8
Table 2: Experimental comparison of state-of-the-art large multi-modal models on our 3DSRBench. Results show that state-of-the-art LMMs exhibit limited 3D spatial reasoning capabilities. Please refer to Sec.˜4.2 for detailed analyses.
@@
Table 4: Experimental results on our 3DSRBench-synthetic-common and 3DSRBench-synthetic-uncommon. We study the robustness of 3D spatial reasoning capabilities of LMMs by analyzing the performance gap between the two splits with images from the same 3D scene but from “common” and “uncommon” viewpoints. We find that LMMs does not generalize well to images with 6D camera viewpoints less represented in their training set. See Sec.˜4.3 for detailed discussions.
We benchmark a wide range of open-sourced and proprietary LMMs on our 3DSRBench-real and analyze 3D spatial reasoning abilities on different types of questions. We consider three baseline results: (i) random: a simple baseline that predicts random answers for all visual questions. (ii) random++: a stronger random baseline that predicts consistent answers given different choice orders of a same visual question in CircularEval. (iii) human: a human-level performance established by human evaluators that did not participate in the data annotation process. We report the full results in Tab.˜2.
@@
We study how design choices of visual encoders can benefit 3D spatial reasoning abilities. Built on LLaVA-v1.5-7B [38], we experiment on a range of models with different choices of visual foundation models, i.e., CLIP [48], MAE [24], DINOv2 [47], SAM [33], or model designs, i.e., mixed encoders and visual projectors. Results in Tab.˜3 show that with mixed encoders, DINOv2 can improve the overall 3D spatial reasoning abilities of LMMs, specifically for orientation and multi-object reasoning questions that build heavily on object 3D orientations. We also notice significant improvements for height questions when adopting MAE and SAM as vision encoder, suggesting that having richer visual features could help localize objects better. With spatial vision aggregator (SVA) [54], we can further improve the LMM with mixed encoder from 37.2% to 37.8%, demonstrating that fusing the semantic features with 3D-aware features from DINOv2 would benefit subsequent reasoning.
@@
We study how the scaling of language model, i.e., in terms of the number of parameters, helps improve the 3D spatial reasoning abilities of LMMs. We consider two series of open-sourced LMMs, QWen2.5 [7] and InternVL2.5 [15], with a range of language model sizes from 0.5B to 72B. From the results in Fig.˜4, we see that the scaling of language model sizes effectively improves the 3D spatial reasoning abilities of LMMs. Larger language models with more parameters exhibit enhanced reasoning abilities. They better capture 3D-aware information from the visual features and perform more complicated 3D spatial reasoning. However, given the importance of 3D spatial reasoning in a broad range of applications, scaling up language model size is highly inefficient — LMMs with over 70B parameters exceed the computation capacity of common robotics or embodied AI systems and significantly limit the model throughput.